# SOUNDCOMPASS: NAVIGATING TARGET SOUND EXTRACTION WITH EFFECTIVE DIRECTIONAL CLUE INTEGRATION IN COMPLEX ACOUSTIC SCENES

### v2

@@ 1 Introduction
Recent studies in deep learning have investigated TSE using auxiliary clues that guide the model toward the target source. Illustrative auxiliary clues encompass class labels [21, 6], text descriptions [14, 13], visual cues [16], or their combinations [11]. In addition to these, the direction of arrival (DoA) has been utilized as a notable clue that leverages the spatial characteristic to isolate a target from interfering sources, irrespective of temporal or spectral attributes.
@@
SPIN input feature. To build a general input feature capturing rich spatial information, we propose a Spectral Pairwise INteraction (SPIN) module that captures all pairwise interactions between sinusoidal components on complex spectrogram of multichannel signals. SPIN features are then fused with a DoA clue in overlapping frequency subbands to capture frequency-dependent spatial cues, extending the advantages of prior band-split approaches [12, 23, 26, 22].
SH embedding for DoA clues. We employ spherical harmonics (SH) as the DoA clue embedding, which provides a continuous angular representation across the 2D sphere. The SH embedding enables the model to handle any DoA value without discretization.
Iterative refinement with temporal clue. Furthermore, we adopt an iterative refinement strategy inspired by the chain-of-inference (CoI) paradigm [24, 8, 10], where the estimated temporal activation is recursively fused with the DoA clue, enabling the model to improve separation quality under challenging multi-source conditions progressively.
@@ 2.2 Directional Clue Integration
Spherical harmonics (SH) embedding. As an accurate and continuous representation of the DoA clue $(\theta,\phi)$, we employ spherical harmonics (SH) as embeddings, as depicted in Fig. 1(b). Unlike one-hot embeddings, SH embeddings allow for the representation of angles without the need for discretization. Additionally, in contrast to cyc-pos embeddings, which define azimuth and elevation angles separately, SH embeddings can describe the position on an $S^{2}$ sphere without coordinate separation, thus providing a consistent representation regardless of coordinate rotation. The complex spherical harmonics of order $n$ and degree $m$ are defined as
@@
where $P_{n}^{m}(\cdot)$ denotes associated Legendre functions. We used up to the 5-th order encoding by stacking the real and imaginary components of spherical harmonics, yielding an embedding vector of dimension $2(N+1)^{2}$ for $N=5$.

### v3

@@ 1 Introduction
Recent studies have investigated TSE using auxiliary clues that guide the model toward the target source, particularly in deep learning-based frameworks. Illustrative auxiliary clues encompass class labels [22, 6], text descriptions [14, 13], visual cues [16], or their combinations [11]. In addition to these, the direction of arrival (DoA) has been utilized as a notable clue that leverages the spatial characteristic to isolate a target from interfering sources, irrespective of temporal or spectral attributes.
@@
Spectral Pairwise INteraction (SPIN) input feature. To build a general input feature capturing rich spatial information, we propose a Spectral Pairwise INteraction (SPIN) module that captures all pairwise interactions between sinusoidal components on complex spectrogram of multichannel signals. SPIN features are then fused with a DoA clue in overlapping frequency subbands to capture frequency-dependent spatial cues, extending the advantages of prior band-split approaches [12, 24, 27, 23].
Spherical harmonics (SH) embedding for DoA clues. We employ spherical harmonics (SH) as the DoA clue embedding, which provides a continuous angular representation across the 2D sphere. The SH embedding enables the model to handle any DoA value without discretization.
Iterative refinement with temporal clue. Furthermore, we adopt an iterative refinement strategy inspired by the chain-of-inference (CoI) paradigm [25, 8, 10], where the estimated temporal activation is recursively fused with the DoA clue to the subsequent stage, enabling the model to improve separation quality under challenging multi-source conditions progressively.
@@ 2.2 Directional Clue Integration
Spherical harmonics (SH) embedding. As an accurate and continuous representation of the DoA clue $(\theta,\phi)$, we employ spherical harmonics (SH) [20] as embeddings, as depicted in Fig. 1(b). Unlike one-hot embeddings, SH embeddings allow for the representation of angles without the need for discretization. Additionally, in contrast to cyc-pos embeddings, which define azimuth and elevation angles separately, SH embeddings can describe the position on an $S^{2}$ sphere without coordinate separation, thus providing a consistent representation regardless of coordinate rotation. The complex spherical harmonics of order $n$ and degree $m$ are defined as
@@
where $P_{n}^{m}(\cdot)$ denotes associated Legendre functions. We use up to the 5-th order encoding by stacking the real and imaginary components of spherical harmonics, yielding an embedding vector of dimension $2(N+1)^{2}$ for $N=5$.
@@ 4 Conclusion
5 Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant (No. RS-2024-00337945), the STEAM research grant (No. RS-2024-00464269) funded by the Ministry of Science and ICT of Korea government (MSIT), Hyundai Motor Company and Kia (No. G01250273), and the BK21 FOUR program through the NRF grant funded by the Ministry of Education of Korea government (MOE).